If a torque of 55.0 N-m is required to turn a bolt and the largest force you can exert is 135 N, how long a lever arm must you use to turn the bolt?


1
Expert's answer
2021-06-21T11:41:10-0400

By the definition of torque, we have:


"\tau =Fr,""r=\dfrac{\tau}{F}=\dfrac{55.0\ N\cdot m}{135\ N}=0.41\ m."